# A Trifecta of Modelling Tools: Navigating COMPAS, Shark and Bayesian Inference for Binary Black Hole Model Selection

### v1

@@ 4 Bayesian inference and model comparison
In this section we outline our procedure for comparing different COMPAS models to data from GWTC-1 (Abbott et al., 2019a), GWTC-2 (Abbott et al., 2021a) and GWTC-3 (Abbott et al., 2023b) including our rescaling methodology. We validate this by method by comparing the likelihoods given the data for the four models from Rauf et al. (2023) for which we have both ‘brute-force’ measurements of the GW rate and our rescaled calculations.
@@ 4.1 Selection criteria
    where $h(f)$ is the waveform measured by the detector, $S(f)$ is the power spectrum density and $f_{\rm min}$ and $f_{\rm max}$ are the minimum and maximum limits for the considered frequency range (Farr et al., 2015; Barrett et al., 2018).
@@
Our supersample of GW events for querying the PDF for each COMPAS model consists of confirmed BBHs and BH-mass-gap events such as GW190521 (Abbott et al., 2020a) and GW190814 (Abbott et al., 2020c). This gives a total of 93 catalog GW detections. Of these, the following GW events are not included in our analysis for the following reasons:
@@ 4.2 Fitting LVK posterior for various COMPAS models
Our basic fitting methodology is based on the widely-used Bayesian framework (e.g., Stevenson et al., 2015; Bouffanais et al., 2019; Mandel et al., 2019; Vitale et al., 2022; Périgois et al., 2023). Given the BBH GW data from some observing runs, $\mathcal{H}$, the posterior for a model specified via hyper-parameters $\lambda$ (for instance a certain mass transfer efficiency $f_{\rm MT}$ or specific angular momentum loss $f_{\gamma}$), takes the form of an inhomogeneous Poisson distribution,
@@
$N_{\lambda}$ is the number of GW events predicted by the model in a given time-frame, while $\mu_{\lambda}$ is the number the model suggests we would actually detect. For convenience we can use these to define some detection efficiency $\beta(\lambda)=\frac{\mu_{\lambda}}{N_{\lambda}}$. $N_{\rm obs}$ is the actual number of detected events in our dataset. We assume that a priori all hyper-parameter combinations are equally likely, so $\pi(\lambda)=1$. Finally, $\mathcal{I}^{k}$ denotes the likelihood of observing GW event $h^{k}$ given the hyper-parameters $\lambda$, written as a marginalisation over the binary properties predicted by that model,
@@
We make model inferences using both of these posteriors in our work and compare how our conclusions change when $N_{\lambda}$ is included as information.
In order to evaluate our posteriors, we first calculate the model volumetric rate using a variant of Eq. 19 but kept in separate metallicity and birth time bins. Given $T_{\rm obs}$, the merger rate and the appropriate volumetric factors, we then find the number of GW events that would fall in each metallicity and time bin, $N_{\lambda,i,j}$, and draw that many events (characterised by a chirp mass and mass ratio) from our full list of $10^{6}$ COMPAS binaries. [^4] We find $N_{\rm COMPAS,i,j}$ in each metallicity and snapshot bins from the COMPAS run. We calculate a weight defined as $N_{\lambda,i,j}/N_{\rm COMPAS,i,j}$ to avoid repeatedly drawing from the COMPAS run. We apply this weight to the PDF. When calculating the detected PDF, we multiply this weight by $p_{\rm det}$. We store the chirp mass, mass ratio and merger redshift of the events across all metallicity and birth time bins before then building the PDF $p_{\rm\textsc{sim}}(\theta|\lambda)$ using Kernel Density Estimation, in particular using the fast KDEpy.FFTKDE package. The bandwidth for our KDE is calculated using Silverman’s rule of thumb (Silverman, 1986) for each hyper-parameter distribution and averaged to obtain the single band-width for the full 3D distribution. As a by-product of method we can also estimate $N_{\lambda}$ and $\mu_{\lambda}$ at the same time by simply keeping track of how many COMPAS binaries we are drawing.
@@
  • The models follow the chirp mass distribution reasonably well except for the chirp mass range 100–130 M. These models predict high mass BHs that either may have not been detected due to the rarity of finding these types of events with current detectors, as they may be associated with noise in the low frequency band, or as a result of inaccuracies in the models resulting in an incorrect representation of the observed population.
@@
  • The shape of the redshift distributions for the models fit the observed distribution quite well. Overall, the Fryer et al. (2012) model is the best fit, while the other models have a wider distribution. This indicates BBHs from these models are more likely to be detected at higher redshifts.
Figure 4: Distribution of $\mathcal{M}_{c}$, $q$ and $z$ for different models. Difference colours refer to the different remnant mass prescriptions investigated in Rauf et al. (2023). The grey shaded regions are the inferred distribution from the LVK collaborations, calculated using the posterior samples for the $N_{\rm obs}=79$ GW events based on the SNR $\geq$ 8 cut.
@@
Figure 5: Normalised histograms of the detected number of BBH merger events calculated using Eq. 27 over the LVK observing time for the last three observing runs. The coloured distributions are obtained from the predictions from the model for 170 sub-samples. The grey band represents the uncertainty in the observed number of BBH mergers depending on the selection criteria adopted (see Section 4.1). The dotted lines are the number of events predicted using the true merger rate from Rauf et al. (2023).
However, if we have only one sample of $N_{\rm sample}$ binaries for a COMPAS, we require an alternative approach to assign an error for these models. [^6] Given the number of models we test in this work, the alternatives of simply using more samples or bootstrapping using the $N_{\rm sample}=10^{6}$ samples would be computationally challenging. Furthermore, subsampling or bootstrapping to generate an error would fail to capture the impact of progenitors of rare events that are not represented in the $N_{\rm sample}=10^{6}$, but which are present in the distributions obtained from our four models with $N_{\rm sample}=170\times 10^{6}$, and from which our errors are extrapolated. We find that the ratio of the standard deviation to the mean number of events as a function of $\log N_{\lambda}$ for the four models has a linear trend such that the relative error increases as the total number of GW events generated by the model decreases, in a roughly Poissonion fashion. In Fig. 6 we show the linear fits for the standard deviation in $N_{\lambda}$ and $\mu_{\lambda}$, which we show in the following equations:
@@
We can now apply this to single samples, assuming similar Gaussian distributions for other COMPAS models. To justify this, we find that the mean $\mu_{\lambda}$ over the 170 sub-samples in Fig. 5 is equivalent to the $\mu_{\lambda}$ we obtain using the true merger rate from the full COMPAS output (one sample) for all four models that we test.
@@ 4.4 Comparing posteriors with and without rescaling merger rate
Figure 8: $\mathcal{I}^{k}$ term for each GW event with $\rm SNR\geq 8$. Each colour represents a different remnant mass prescription from COMPAS and the error bar represents the sampling error over 170 samples when building the simulated PDF of $\mathcal{M}_{c}$, $q$ and $z$ given the model.
Fig. 8 shows the $I^{k}$ term for individual GW events. This is a practical approach to determine the probability of individual GW events occurring in the populations generated by the COMPAS models. The Schneider et al. (2021) model generally returns high probabilities when querying the PDF, while the Fryer et al. (2012) model is disfavoured. We find some events are outliers and are not well reproduced by some or all of the models relative to the other events. As an example, Mandel & Müller (2020) with $f_{\rm WR}=0.2$ and Schneider et al. (2021) are less likely to output the progenitors of GW190814, GW190917_114630 and GW200210_092254 compared to Fryer et al. (2012) and Mandel & Müller (2020). For GW190426_190642, GW190602_175927 and GW191109_010717 Schneider et al. (2021) is distinctly preferred over the other models, although in these cases, and especially that of GW190521, the models are generally less able to produce the required event-level properties. This is expected for GW190521 as both masses, $\sim 85M_{\odot}$ and $\sim 66M_{\odot}$, most likely lie in the pair-instability supernovae mass gap, making it atypical of binaries formed via the isolated formation channel (Abbott et al., 2020a).
Figure 9: Normalised histograms of the log likelihoods obtained using 170 sub-samples for the models explored in Rauf et al. (2023) with and without the merger rate dependence. The top two panels are the likelihoods comparing the models to the GW events that fit the $\rm FAR\leq 0.25yr^{-1}$ and $p_{\rm astro}\geq 0.9$ criteria, while the bottom two panels compares to the $\rm SNR\geq 8$ criteria. The dotted lines are the likelihoods using the true merger rates calculated from Rauf et al. (2023). In all cases we see that the true likelihood is well represented within our distribution of 170 subsamples, indicating our use of a relatively smaller number of $10^{6}$ binaries is not expected to bias our results.
@@ 5 Results
Having validated our methodology using the four models analysed in Rauf et al. (2023), this section discusses the outputs from our new set of COMPAS models where we vary the range of hyper-parameters described previously in Section 2 (see also Table 1. We start with describing the trends in the number of detected events when varying the COMPAS hyper-parameters. We calculate the Bayes factor to identify groups of models that are more preferred. We then show our likelihood analysis, where we discuss the best and worst models when compared to the observed population parameters. We follow up this analysis and look at specific events and their $\mathcal{I}^{k}$ value to determine which models are more and less likely to generate the progenitors of these events. See Tables 4, 5 and 6 in Appendix C for a detailed description of the hyper-parameters for all 304 COMPAS models in this analysis.
@@ 5.1 $\mu_{\rm GW}$ versus COMPAS parameters
Figure 10: Detected number of BBH merger events versus the mass transfer fraction for no CHE at fixed $f_{\gamma}$. Left: We fix the common envelope evolution parameter $\alpha_{\rm CE}=0.1$ and show results for different specific angular momentum loss parameters. Right: We show the models for any $f_{\gamma}$ and $\alpha_{\rm CE}$ which output $\mu_{\lambda}$ close to the number of observed LVK events given by the grey band.
@@
In the left plot of Fig. 10, we see $\mu_{\rm GW}$ increase with mass transfer fraction, $f_{\rm MT}$, when the corresponding angular momentum is lost directly from the accretor ($f_{\gamma}=0$). Increasing the mass transfer can lead to more equal mass objects, which we find most of our COMPAS models producing. This trend disappears for higher $f_{\gamma}$. We credit this to non-trivial evolutionary effects such as the stability prescriptions, that determine any future interactions within the binary. This preference for intermediate values of $f_{\gamma}$ agrees with the findings by Santoliquido et al. (2021). Irrespective of $f_{\rm MT}$, $\mu_{\rm GW}$ peaks for moderate values of $f_{\gamma}=0.5$ but drops again for high $f_{\gamma}$ when angular momentum loss occurs closer to the L2 point. At this point, mass is lost at higher velocities so more angular momentum is lost from the system, which in turn can induce CE evolution and merger during the mass transfer stage, stopping the formation of a BBH. These results are in agreement with Lu et al. (2023); van Son et al. (2022b) that the number of stable mass channel objects increases for decreasing $f_{\gamma}$, as opposed to the objects formed by CE evolution. Marchant et al. (2021) also found that the stable channel dominates the BBH merger rate. We note that when $f_{\rm MT}=1$, $f_{\gamma}$ should have no impact. Any variations in the number of events at $f_{\rm MT}=1$ is hence due to Eddington-limited accretion occurring after the first mass-transfer event(s), once the accretor has become a BH.  [^7] As a result of this finding, dual functionality in the AM loss was later implemented in COMPAS, as discussed in Section 2.4. If specific AM loss is set at the compact object as well as prior to the formation of the first compact object we expect the number of events will no longer vary for $f_{\rm MT}=1$, although at time of writing this functionality is yet to be comprehensively tested.
We show the models that produce the closest $\mu_{\lambda}$ to the observed LVK events in the right plot of Fig. 10. We find that for $\mu_{\rm GW}$ to be below 100 events (closer to the number of detected events from all observing runs), a CE efficiency $\alpha_{\rm CE}=10,100$ and $f_{\gamma}=0.75,1.0$ with no CHE is required. This implies that the CE is ejected earlier in the evolution and the orbit tightens at a wider separation, which means longer merger times and less events in the same observed time. This result validates the studies by Santoliquido et al. (2021), Bouffanais et al. (2021) and Grichener (2023). With the exception of these models, the CE efficiency has no significant impact on the populations, as the range for the number of events does not vary significantly with $\alpha_{\rm CE}$.

### v2

@@ 4 Bayesian inference and model comparison
In this section we outline our procedure for comparing different COMPAS models to data from GWTC-1 (Abbott et al., 2019a), GWTC-2 (Abbott et al., 2021a) and GWTC-3 (Abbott et al., 2023b) including our rescaling methodology. We validate this method by comparing the likelihoods given the data for the four models from Rauf et al. (2023) for which we have both ‘brute-force’ measurements of the GW rate and our rescaled calculations.
@@ 4.1 Selection criteria
    where $h(f)$ is the waveform measured by the detector, $h^{*}(f)$ is its Hermitian conjugate, $S(f)$ is the power spectrum density, $\textrm{Re}[]$ is the real component of the inner product, and $f_{\rm min}$ and $f_{\rm max}$ are the minimum and maximum limits for the considered frequency range (Farr et al., 2015; Barrett et al., 2018).
@@
Our supersample of GW events for querying the probability distribution function (PDF) for a set of parameters in each COMPAS model (see Section 4.2) consists of confirmed BBHs and BH-mass-gap events such as GW190521 (Abbott et al., 2020a) and GW190814 (Abbott et al., 2020c). This gives a total of 93 catalog GW detections. Of these, the following GW events are not included in our analysis for the following reasons:
@@ 4.2 Fitting LVK posterior for various COMPAS models
Our basic fitting methodology is based on the widely-used Bayesian framework (e.g., Stevenson et al., 2015; Bouffanais et al., 2019; Mandel et al., 2019; Vitale et al., 2022; Périgois et al., 2023). Given the BBH GW data from some observing runs, $\mathcal{H}$, the posterior for a model specified via hyper-parameters $\lambda$ (for instance a certain mass transfer efficiency $f_{\rm MT}$ or specific AM loss $f_{\gamma}$), takes the form of an inhomogeneous Poisson distribution,
@@
$N_{\lambda}$ is the number of GW events predicted by the model in a given time-frame, while $\mu_{\lambda}$ is the number the model suggests we would actually detect. $N_{\rm obs}$ is the actual number of detected events in our dataset. We assume that a priori all hyper-parameter combinations are equally likely, so $\pi(\lambda)=1$. Finally, $\mathcal{I}^{k}$ denotes the likelihood of observing GW event $h^{k}$ given the hyper-parameters $\lambda$, written as a marginalisation over the binary properties predicted by that model,
@@
For convenience we can use $N_{\lambda}$ and $\mu_{\lambda}$ to define some detection efficiency $\beta(\lambda)=\frac{\mu_{\lambda}}{N_{\lambda}}$. We make model inferences using both of these posteriors in our work and compare how our conclusions change when $N_{\lambda}$ is included as information.
In order to evaluate our posteriors, we first calculate the model volumetric rate using a variant of Eq. 19 but kept in separate metallicity and birth time bins. Given $T_{\rm obs}$, the merger rate and the appropriate volumetric factors, we then find the number of GW events that would fall in each metallicity and time bin, $N_{\lambda,i,j}$, and draw that many events (characterised by a chirp mass and mass ratio) from our full list of $10^{6}$ COMPAS binaries. [^8] We find $N_{\rm COMPAS,i,j}$ in each metallicity and snapshot bins from the COMPAS run. We calculate a weight defined as $N_{\lambda,i,j}/N_{\rm COMPAS,i,j}$ to avoid repeatedly drawing from the COMPAS run. We apply this weight to the PDF. When calculating the detected PDF, we multiply this weight by $p_{\rm det}$. We store the chirp mass, mass ratio and merger redshift of the events across all metallicity and birth time bins before then building the PDF $p_{\rm\textsc{sim}}(\theta|\lambda)$ using Kernel Density Estimation (KDE), in particular using the fast KDEpy.FFTKDE package. The bandwidth for our KDE is calculated using Silverman’s rule of thumb (Silverman, 1986) for each hyper-parameter distribution and averaged to obtain the single band-width for the full 3D distribution. As a by-product of method we can also estimate $N_{\lambda}$ and $\mu_{\lambda}$ at the same time by simply keeping track of how many COMPAS binaries we are drawing.
@@
  • With the exception of Schneider et al. (2021), the models replicate the first peak of the observed chirp mass distribution well. No models fit the second peak around $30M_{\odot}$. For all models, there is zero probability of producing chirp masses $>40M_{\odot}$ until 100–130 M. This implies these models predict high mass BHs that either may have not been detected due to the rarity of finding these types of events with current detectors, as they may be associated with noise in the low frequency band, or as a result of inaccuracies in the models resulting in an incorrect representation of the observed population.
@@
  • The shape of the redshift distributions for the models fit the observed distribution quite well. Overall, the Fryer et al. (2012) model is the best fit, while the other models have a wider distribution. This indicates BBHs from these models are more likely to be detected at higher redshifts (closer to $z\sim 1$) compared to Fryer et al. (2012).
Figure 4: Distribution of chirp mass, mass ratio and redshift for different models. Different colours refer to the different remnant mass prescriptions investigated in Rauf et al. (2023). The grey shaded regions are the inferred distribution from the LVK collaborations, calculated using the posterior samples for the $N_{\rm obs}=79$ GW events based on the SNR $\geq$ 8 cut.
@@
Figure 5: Normalised histograms of the detected number of BBH merger events calculated using Eq. 27 over the LVK observing time for the last three observing runs. The coloured distributions are obtained from the predictions from the model for 170 sub-samples. The grey band represents the uncertainty in the observed number of BBH mergers depending on the selection criteria adopted (see Section 4.1). The dotted lines are the number of events predicted using the Rauf et al. (2023) merger rate from Rauf et al. (2023).
However, if we have only one sample of $N_{\rm sample}$ binaries for a COMPAS output, we require an alternative approach to assign an error for these models. Given the number of models we test in this work, the alternatives of simply using more samples or bootstrapping using the $N_{\rm sample}=10^{6}$ samples would be computationally challenging. Furthermore, subsampling or bootstrapping to generate an error would fail to capture the impact of progenitors of rare events that are not represented in the $N_{\rm sample}=10^{6}$, but which are present in the distributions obtained from our four models with $N_{\rm sample}=170\times 10^{6}$, and from which our errors are extrapolated. We find that the ratio of the standard deviation to the mean number of events as a function of $\log N_{\lambda}$ for the four models has a linear trend such that the relative error increases as the total number of GW events generated by the model decreases, in a roughly Poissonion fashion. In Fig. 6 we show the linear fits for the standard deviation in $N_{\lambda}$ and $\mu_{\lambda}$, which we show in the following equations:
@@
We can now apply this to single samples, assuming similar Gaussian distributions for other COMPAS models. To justify this, we find that the mean $\mu_{\lambda}$ over the 170 sub-samples in Fig. 5 is equivalent to the $\mu_{\lambda}$ we obtain using the Rauf et al. (2023) merger rate from the full COMPAS output (one sample) for all four models that we test.
@@ 4.4 Comparing posteriors with and without rescaling merger rate
Figure 8: The likelihood, $\mathcal{I}^{k}$, for each GW event with $\rm SNR\geq 8$. Each colour represents a different remnant mass prescription from COMPAS and the error bar represents the sampling error over 170 samples when building the simulated PDF of $\mathcal{M}_{c}$, $q$ and $z$ given the model.
Fig. 8 shows the $\mathcal{I}^{k}$ term for individual GW events. This is a practical approach to determine the probability of individual GW events occurring in the populations generated by the COMPAS models. The Schneider et al. (2021) model generally returns high probabilities when querying the PDF, while the Fryer et al. (2012) model is disfavoured. We find some events are outliers and are not well reproduced by some or all of the models relative to the other events. As an example, Mandel & Müller (2020) with $f_{\rm WR}=0.2$ and Schneider et al. (2021) are less likely to output the progenitors of GW190814, GW190917_114630 and GW200210_092254 compared to Fryer et al. (2012) and Mandel & Müller (2020). For GW190426_190642, GW190602_175927 and GW191109_010717 Schneider et al. (2021) is distinctly preferred over the other models, although in these cases, and especially that of GW190521, the models are generally less able to produce the required event-level properties. This is expected for GW190521 as both masses, $\sim 85M_{\odot}$ and $\sim 66M_{\odot}$, most likely lie in the pair-instability supernovae mass gap, making it atypical of binaries formed via the isolated formation channel (Abbott et al., 2020a).
Figure 9: Normalised histograms of the log likelihoods obtained using 170 sub-samples for the models explored in Rauf et al. (2023) with and without the merger rate dependence. The top two panels are the likelihoods comparing the models to the GW events that fit the $\rm FAR\leq 0.25yr^{-1}$ and $p_{\rm astro}\geq 0.9$ criteria, while the bottom two panels compares to the $\rm SNR\geq 8$ criteria. The dotted lines are the likelihoods using the Rauf et al. (2023) merger rate. In all cases we see that the true likelihood is well represented within our distribution of 170 subsamples, indicating our use of a relatively smaller number of $10^{6}$ binaries is not expected to bias our results.
@@ 5 Results
Having validated our methodology using the four models analysed in Rauf et al. (2023), this section discusses the outputs from our new set of COMPAS models where we vary the range of hyper-parameters described previously in Section 2 (see also Table 1). We start with describing the trends in the number of detected events when varying the COMPAS hyper-parameters. We calculate the Bayes factor to identify groups of models that are more preferred. We then show our likelihood analysis, where we discuss the best and worst models when compared to the observed population parameters. We follow up this analysis and look at specific events and their $\mathcal{I}^{k}$ value to determine which models are more and less likely to generate the progenitors of these events. See Tables 4, 5 and 6 in Appendix C for a detailed description of the hyper-parameters for all 304 COMPAS models in this analysis.
@@ 5.1 $\mu_{\rm GW}$ versus COMPAS parameters
Figure 10: Detected number of BBH merger events versus the mass transfer fraction for no CHE at fixed $f_{\gamma}$. Left: We fix the CE evolution parameter $\alpha_{\rm CE}=0.1$ and show results for different specific AM loss parameters. Right: We show the models for any $f_{\gamma}$ and $\alpha_{\rm CE}$ which output $\mu_{\lambda}$ close to the number of observed LVK events given by the grey band.
@@
In the left plot of Fig. 10, we see $\mu_{\rm GW}$ increase with mass transfer fraction, $f_{\rm MT}$, when the corresponding AM is lost directly from the accretor ($f_{\gamma}=0$). Increasing the mass transfer can lead to more equal mass objects, which we find most of our COMPAS models producing. This trend disappears for higher $f_{\gamma}$. We credit this to non-trivial evolutionary effects such as the stability prescriptions, that determine any future interactions within the binary. This preference for intermediate values of $f_{\gamma}$ agrees with the findings by Santoliquido et al. (2021). Irrespective of $f_{\rm MT}$, $\mu_{\rm GW}$ peaks for moderate values of $f_{\gamma}=0.5$ but drops again for high $f_{\gamma}$ when AM loss occurs closer to the L2 point. At this point, mass is lost at higher velocities so more AM is lost from the system, which in turn can induce CE evolution and merger during the mass transfer stage, stopping the formation of a BBH. These results are in agreement with Lu et al. (2023); van Son et al. (2022b) that the number of stable mass channel objects increases for decreasing $f_{\gamma}$, as opposed to the objects formed by CE evolution. Marchant et al. (2021) also find that the stable channel dominates the BBH merger rate. We note that when $f_{\rm MT}=1$, $f_{\gamma}$ should have no impact. Any variations in the number of events at $f_{\rm MT}=1$ is hence due to Eddington-limited accretion occurring after the first mass-transfer event(s), once the accretor has become a BH.  [^10] As a result of this finding, dual functionality in the AM loss was later implemented in COMPAS, as discussed in Section 2.4. If specific AM loss is set at the compact object as well as prior to the formation of the first compact object we expect the number of events will no longer vary for $f_{\rm MT}=1$, although at time of writing this functionality is yet to be comprehensively tested.
We show the models that produce the closest $\mu_{\lambda}$ to the observed LVK events in the right plot of Fig. 10. We find that for $\mu_{\rm GW}$ to be below 100 events (closer to the number of detected events from all observing runs), a CE efficiency $\alpha_{\rm CE}=10,100$ and $f_{\gamma}=0.75,1.0$ with no CHE is required. This implies that the CE is ejected earlier in the evolution and the orbit tightens at a wider separation, which means longer merger times and less events in the same observed time. This result validates the studies by Santoliquido et al. (2021), Bouffanais et al. (2021) and Grichener (2023). As discussed, $\alpha_{\rm CE}=100$ is extreme, which suggests an alternative approach for the CE is required for forming populations with a comparable number of GW events to LVK. With the exception of these models, the CE efficiency has no significant impact on the populations, as the range for the number of events does not vary significantly with $\alpha_{\rm CE}$.